OpenStudy (anonymous):

Determine the number of degrees that the central angle must measure for a circle graph if it is to show that a person slept 8 hours of the day.

OpenStudy (anonymous):

8° 24° 40° 120° 160°

OpenStudy (anonymous):

sorry. didn't mean to put thaqt.

OpenStudy (anonymous):

*that (I can't type)

OpenStudy (anonymous):

360/8=45

OpenStudy (anonymous):

45 isn't an option :/

OpenStudy (anonymous):

120?

OpenStudy (anonymous):

That is an option!

OpenStudy (anonymous):

i am not sure though..

OpenStudy (anonymous):

Me either. :/

OpenStudy (anonymous):

my 60% ;)

OpenStudy (anonymous):

I will guess and let you know later!

OpenStudy (anonymous):

sure..

OpenStudy (amistre64):

8 out of 24 8/24 = 1/3 360/3 = 120

OpenStudy (anonymous):

i got it right.. lol